How many gram does nitrogen (N) present in 100 Kg ammonium nitrate ((NH4NO3) fertilizer?


1
Expert's answer
2021-04-06T06:06:04-0400

M (NH4NO3) = 80.04 g/mol

n (NH4NO3) = m / M = 100 000 / 80.04 = 1249.33 mol

The ratio of moles of N to moles of NH4NO3 is 2:1. Therefore, N moles is:

n (N) = 1249.33 / 1 x 2 = 2498.66 mol

M (N) = 14.01 g/mol

m (N) = n x M = 2498.66 x 14.01 = 35006.18 g = 35 kg